### v6

# IntelliRadar: A Comprehensive Platform to Pinpoint Malicious Packages Information from Cyber Intelligence

Abstract.
Malicious packages in public registries pose serious threats to software supply chain security. While current software component analysis (SCA) tools rely on databases like OSV and Snyk to detect these threats, these databases suffer from delayed updates and incomplete coverage. However, they miss intelligence from unstructured sources like social media and developer forums, where new threats are often first reported. This delay extends the lifecycle of malicious packages and increases risks for downstream users.
To address this, we developed a novel and comprehensive approach to construct a platform IntelliRadar to collect disclosed malicious package names from unstructured web content. Specifically, by exhaustively searching and snowballing the public sources of malicious package names, and incorporating large language models (LLMs) with domain-specialized Least to Most prompts, IntelliRadar ensures comprehensive collection of historical and current disclosed malicious package names from diverse unstructured sources. As a result, we constructed a comprehensive malicious package database containing 34,313 malicious NPM and PyPI package names. Our evaluation shows that IntelliRadar achieves high performance (97.91% precision) on malicious package intelligence extraction. Compared to existing databases, IntelliRadar identifies 7,542 more malicious package names than OSV and 12,684 more than Snyk. Furthermore, 76.6% of NPM components and 70.3% of PyPI components in IntelliRadar were collected earlier than in Snyk’s database. IntelliRadar is also more cost-efficient, with a cost of \$0.003 per piece of malicious package intelligence and only \$7 per month for continuous monitoring. Furthermore, we identified and received confirmation for 1,981 malicious packages in downstream package manager mirror registries through the IntelliRadar.
Supply Chain Security, Malicious Packages, Cyber Threat Intelligence, Third-Party Library
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773113isbn: 979-8-4007-2025-3/2026/04ccs: Software and its engineering Software libraries and repositories
1. Introduction
With the widespread global use of open-source software, malicious actors have discovered an effective means to disseminate malicious code through open-source platforms. Especially, TPL registries, such as NPM and PyPI, have become disaster areas of malicious code, in which attackers deliberately upload packages, embedded with malicious code, and induce downstream users to include them as dependencies. These malicious packages often carry viruses, Trojans, ransomware, etc. (npm_remote_trojans; npm_stole_pwd; pypi_malicious_trojans), and stealthily infiltrate user systems by masquerading as regular libraries or tools. Unlike unintentional vulnerabilities, these are deliberately created by attackers with malicious intent.
To address the threat of malicious packages, academic and industrial researchers have worked to prevent their spread by investigating malicious package types, taxonomies (guo2023empirical; zhou2024large; zhou2024oss), and attack surfaces (okafor2022sok; ladisa2023journey; ladisa2023sok), as well as developing detection tools (huang2024donapi; li2023malwukong; vu2023bad; zahan2024shifting); simultaneously, software composition analysis tools like Snyk (snykweb), BlackDuck (BlackDuckweb), OWASP Dependency Check (OWASPweb), and Dependabot (dependabotweb) with security databases are used to identify malicious dependencies, relying on platforms such as GitHub Advisory (advisory_database_ref), NVD (nvdweb), and OSV (osv_database_ref) for updates. However, despite the abundance of detection methods, the malicious packages discovered by these researchers are typically disclosed through scattered unstructured web pages, making it difficult for this information to effectively reach downstream users and developers. Consequently, many identified malicious packages still remain on mirror servers. Up to 72.4% of malicious PyPI packages remain on mirror servers after being reported (guo2023empirical), with the colorwed package (snyk_colorwed) being downloaded over 1,000 times after identification as malicious, revealing both a lack of security awareness and increased risk as these packages become known to potential attackers during this period.
Therefore, we aim to mitigate these information delays for malicious packages in this paper. Specifically, as detailed by the motivating example in Section 2, the newly identified malicious packages from researchers, companies, or organizations are usually reported on various public channels, such as personal blogs, tweets, or news on reputed platforms. To address this gap, we aim to establish a comprehensive and public-available intelligence platform to automatically collect, process, and identify malicious packages in time.
Figure 1. The Timeline of Intelligence Reporting of the Malicious Package colorwed
To achieve this, we face the following challenges: 1) Intelligence Sources. To ensure the intelligence timeliness, we should timely capture first-hand intelligence when they are posted. Therefore, a comprehensive list of intelligence sources should be monitored. 2) Key Information Extraction. Since different reporters formalize intelligence in various ways, it is still difficult to accurately retrieve the key information of malicious packages, such as package names, corresponding versions, and additional key information (i.e., types and behavior patterns) from intelligence. 3) Trustworthy of Intelligence. The intelligence publicly reported online could also be unreliable. Thus, it is also non-trivial to accurately identify inaccurate information before it reaches downstream users.
To fill these gaps, we propose a comprehensive approach IntelliRadar to construct a platform to pinpoint the coverage, timeliness, and accuracy of automated collection and processing of malicious package intelligence. Specifically, for challenge 1), starting from existing malicious package reports, we conducted a thorough exploration of intelligence sources by summarizing domain-specific keywords and snowballing results from search engines, to identify as many intelligence sources as possible. For challenge 2), we incorporate the large language models (LLMs) to extract the key information of malicious packages precisely. Unlike general cyber threat intelligence approaches such as CTIKG (huang2024ctikg) and SecBERT (host2023constructing), our task requires distinguishing malicious package names from benign packages in the same text and performing ecosystem-specific semantic understanding. To handle potential LLM inaccuracy (i.e., hallucination (huang2023surveyhallucinationlargelanguage)), we introduced a multi-faceted approach to prompt LLMs with domain knowledge of malicious packages. For challenge 3), we conduct an empirical analysis on conflicted information among intelligence, and based on their correctness, we introduce a voting mechanism based on recency to cross-validate intelligence from different sources, for the same malicious packages.
Our experiment and analysis show that 1) Our method collected over 34,313 malicious package names from 24 sources, achieving a clearly higher coverage of malicious packages compared to existing databases (i.e., 12,648 and 7,542 package names are identified as missing in the Snyk Database and OSV database, respectively). 2) Our approach demonstrates exceptional effectiveness in processing and extracting critical information from online threat intelligence sources, achieving 96.4% recall in keyword-based text filtering. By introducing CoT reasoning and few-shot techniques, our model attains an F1 of 94.87%, surpassing other LLMs and prompting-based methods. 3) Our intelligence platform can effectively discover the intelligence of malicious packages earlier than existing platforms. More than 76.6% of NPM and 70.3% of PyPI malicious package names were discovered earlier than Snyk, with 4,711 malicious packages discovered earlier than OSV. 4) Our approach is highly cost-efficient: text filtering reduces consumption by up to 58.4%, and with a monthly cost of \$7, each malicious package intelligence extraction costs only \$0.003 via LLM-based analysis. 5) GitHub and Phylum serve as the primary intelligence sources, documenting 45.85% of NPM and 34.8% of PyPI intelligence, respectively. GitHub, Phylum, Sonatype, OSV, Checkmarx, and Medium collectively account for more than 70% of intelligence in both ecosystems.
We summarize the main contributions as follows:
$\bullet$ We proposed IntelliRadar, a comprehensive LLM-based SSC intelligence analysis platform for the complete and in-time collection of malicious package intelligence, achieving an F1-score of 94.87%.
$\bullet$ We constructed a comprehensive and human-validated dataset containing intelligence on 34,313 malicious package names, establishing the largest database for PyPI and NPM package managers to date, which is publicly accessible through our website (website).
$\bullet$ Our approach demonstrates excellent cost-efficiency, with IntelliRadar requiring only \$7 monthly for monitoring all relevant web pages, and each piece of intelligence costing \$0.003 to identify.
$\bullet$ We reported intelligence on over 1,981 malicious packages to downstream mirror maintainers, significantly contributing to the security of the open-source ecosystem.
Our research adheres to the following ethical principles: (a) We strictly follow website terms of service and robots.txt protocols during data collection; (b) We only collect and process publicly available information, with no attempt to access restricted data.
2. Motivation
Through our analysis of existing malicious package intelligence platforms, we discovered two critical challenges in the current malicious package intelligence ecosystem: (1) delays in threat information propagation and (2) incompleteness of authoritative databases.
$\bullet$ Delay in Intelligence Propagation In the propagation chain of malicious packages, significant delays exist between initial discovery and eventual inclusion in SSC security databases. Research by Jacobs et al.(10.1093/cybsec/tyaa015) indicates that attackers can exploit such information propagation delays to compromise downstream users’ software security. An typical motivating example is illustrated in Figure 1, the propagation timeline of the malicious PyPI package colorwed demonstrates this issue: while the package was discovered and reported on social media(colorwed_twitter) on the same day it was uploaded to the PyPI registry (December 23, 2022), and subsequently reported by multiple security companies (JFrog on November 24(colorwed_jfrog), Phylum and Sontype on December 15 (colorwed_phylum; colorwed_sonatype)), it took approximately one month (December 21) before being included in software composition analysis tools like Snyk’s database. This case clearly demonstrates the critical timing deficiencies in current malicious package intelligence propagation mechanisms.
$\bullet$ Incompleteness of Authoritative Databases Currently, malicious package intelligence across mainstream SSC security platforms exhibits significant limitations. While some platforms (e.g., Snyk, OSV, and GitHub Advisory) provide structured databases of malicious packages, their coverage of the entire ecosystem’s malicious packages remains limited. Concurrently, a substantial volume of malicious package intelligence is disseminated through unstructured formats such as blog posts, security reports, and technical articles. These unstructured sources often contain rich, detailed information about malicious packages, including their behaviors. However, the unstructured nature of this information impedes automated processing. LLMs offer a promising solution to this challenge by extracting valuable intelligence from these natural language sources through their contextual understanding capabilities.
3. Approach
Figure 2. Workflow of the IntelliRadar
We developed a framework named IntelliRadar for collecting, processing, and aggregating open-source malicious package intelligence. As shown in Figure 2, our method contains four parts. 1) IntelliRadar first identify and collect intelligence sources that are highly relevant to the exposure of packages by greedy snowballing searching, after that, 2) it collects all posted intelligences from each source, and extract webpages that possibly contain malicious package related information. Then, 3) IntelliRadar employs LLMs with well-designed Chain of Thoughts (CoT) to precisely extract relevant entities and relationships, and based on them, 4) IntelliRadar further aggregates these relevant details from different intelligences and derive the complete malicious intelligence database.
3.1. Intelligence Source Identification
Identification of the intelligence source is fundamental for intelligence collection. To ensure the completeness of intelligence collection, we first designed a recursive process to identify open source malicious package intelligence sources by 3 major steps : ① keyword selection, ② intelligence source identification, and ③ recursive expansion of intelligence sources. Specifically, as presented in Figure 2, we first construct a comprehensive keyword set to search malicious package related intelligence on the open Internet, and based on the results, we sort out the sources (e.g., websites, accounts, and forums, etc) that constantly post malicious package information, as the targeted package intelligence sources. Moreover, to further ensure that no major sources, especially those post first-hand intelligences, missed, we further introduce a greedy strategy to recursively identify possible intelligence sources documented in identified intelligences, till no new intelligence sources emerges.
3.1.1. Keyword Selection.
We first prepare the keywords for searching of existing malicious package related intelligence on the open Internet. To ensure the coverage and diversity of searched intelligence, we follow a rigorous strategy to extract keywords that are commonly used in existing malicious package reports.
To acquire the existing malicious package reports, we first collect the malicious package list (including 2,351 PyPI and 1,984 NPM packages) from the existing well-known dataset of malicious packages (ohm2020backstabber), and cross-map it with Snyk and OSV databases (osv_database_ref; snyk_database_ref), which are well verified with reference links, to collect all reference intelligence reports. Based on this corpus, we then select the most relevant keywords that are commonly used in these intelligence reports. Specifically, we collect two groups of keywords,
$\bullet$ Common Keywords, we aim to identify the keywords that are most commonly used in these malicious intelligence reports. To this end, we apply the Latent Dirichlet Allocation (LDA) model (tong2016text) to cluster the text topics from the existing reports. After excluding stopwords and irrelevant common words, we selected 17 common keywords that are related to malicious packages out of the top 100 clustered topics.
$\bullet$ Special Keywords, apart from these common keywords, there could also be certain words that are only common in specific report. To this end, we adopt the Term Frequency-Inverse Document Frequency (TF-IDF) algorithm (qaiser2018text) to identify the top 10 most frequent keywords that are not included in the common keywords in each report. After excluding irrelevant words, we aggregate these words as the special keywords. Interestingly, most of these selected special keywords are the names of malicious packages and their specific attributes, such as execution and control. The detailed list of keywords are presented in Table 1.
Table 1. Common and Special Keywords
Type Keywords
Special Keywords ¡Package name¿, ¡Specific attributes¿
Common Keywords package, security, malicious, attacker, account, user, registry, code, software, github, malware, repository, dataset, infected, script, vulnerability, workflow
3.1.2. Intelligence Source Identification.
Based on these keywords, we then conduct a comprehensive searching for malicious package intelligence on the open Internet.
Specifically, we conduct targeted searches through the Google Custom Search API with special keywords configured as mandatory conditions and common keywords configured as optional parameters. Considering that 99.3% of searches yielding no more than 100 records, for each search, we only take the first 100 records as its result. After deduplication, these pages yielded 7,330 unique webpage links spanning 2,412 distinct web domains. These web domains are considered to be possible sources of malicious intelligence sources.
Since sources with only few historical reports of malicious packages are not necessarily to monitor, we further filter out the web domains that are with low frequencies in our search results. Specifically, Our statistics revealed that 80.2% of web domains have only one posts, and only 228 web domains (9.5%) had ever published over 10 posts in our search results. After we manually inspected the content of these posts, we found that only 24 web domains have published malicious package related information in their historical posts, and they are ultimately identified as the intelligence sources that specialize in SSC security and malicious software analysis in this step for further collection and monitoring.
3.1.3. Recursive Expansion of Intelligence Sources.
Moreover, to ensure that our search did not miss necessary intelligence sources, we then conduct a recursive analysis on the identified intelligence sources, aiming to explore new intelligence sources, till no new source emerges. Specifically, for each of the 24 intelligence sources, we extracted all external links from the web contents we collected in the Google search results. After inspecting the 10,326 newly identified external links, we found that 97% of these links are actually from the 24 existing intelligence sources that we have already identified, and the remaining 3% are irrelevant advertising links.
Figure 3. Sources and classification of intelligence sources
3.2. Intelligence Collection
After identifying intelligence sources, we collected all webpage content for historical intelligence extraction and continuous monitoring. As shown in Figure 3, we classified these sources into five functional categories. Each category employs distinct structures for SSC intelligence distribution: Cybersecurity companies utilize specialized second- or third-level domains, Security News sites implement keyword indexing systems, while developer communities, code platforms, and social networks distribute intelligence through specific accounts. These structural patterns enabled targeted collection from relevant subdomains and accounts, eliminating the need for comprehensive site-wide crawling. However, even within these targeted SSC-related intelligence sources, many webpages still contain content irrelevant to malicious package intelligence. To overcome this challenge, we established clear inclusion and exclusion criteria for filtering relevant webpages. Our filtering employs a two-tier approach: ecosystem-specific keywords (e.g., ”pypi,” ”npm”) as mandatory criteria, and common security keywords from  Table 1 (e.g., ”malicious,” ”security,” ”package”) as optional filters. This mechanism significantly reduces computational overhead while maintaining comprehensive coverage. To collect webpage content, we developed specialized HTML parsers targeting information-rich elements (tables, lists, paragraphs, code snippets, headings) and iframes containing malicious package details and IOCs.
3.3. Malicious Package Intelligence Extraction
Figure 4. Entity and Relationship Analysis Using the CoT Prompts
As shown in Figure 4 (1), the web content often includes a significant amount of dispersed information related to malicious packages, which is highlighted in yellow for clarity and emphasis. To efficiently extract the key intelligence from these texts, we employed LLM to enhance the information extraction process (min2023recent; wang2023gpt). However, LLM still faces issues like hallucination, limiting extraction accuracy (ling2023improving). To address this, we first preprocess the web content to extract potential names of malicious packages. Then, we feed both the extracted potential malicious package information and the web content together into LLM to extract more detailed and accurate malicious package information (martino2023knowledge).
The potential malicious package name extraction process is illustrated in Figure 4 (2). We have observed that malicious packages on PyPI and NPM often try to deceive downstream users into downloading them by creating package names that are similar to benign packages, such as through typosquatting (tahir2018s; vu2020typosquatting) creating package names with minor letter variations. As a result, the names of these malicious packages often aren’t real words found in English dictionaries. Based on this observation, we designed a two-step extraction pipeline. First, we employ the regular expression r”(?:@[a-zA-Z0-9-]+/)?[a-zA-Z0-9][a-zA-Z0-9._-]+” to extract strings matching package naming conventions from webpage content. Subsequently, we apply filtering operations to the extracted candidate package names: removing duplicates to eliminate redundancy, stripping trailing punctuation marks such as periods and commas, and filtering out common English vocabulary and stopwords, thereby retaining potential malicious package names.
To analyze malicious package intelligence, we employ CoT reasoning with three components: entity extraction, relationship analysis, and verification. This decomposition aligns with how security analysts process intelligence: extracting discrete data, revealing patterns and connections, then ensuring integrity.
(1) Entity Extraction (Fig 4a): Identify and extract key entities related to malicious packages. The prompt design consists of six components: 1) task description, 2) intelligence content, 3) 9 entity types including Package Name, Version, Date of Discovery, Repository URL, Method of Attack, Discoverer, Impacted Systems, Attack Vector, and IOCs, 4) potential package names from previous steps, 5) common malicious naming patterns (typosquatting and misspelling), and 6) few-shot examples. The LLM employs a Chain-of-Thought approach with progressive difficulty: first confirming malicious package names, then extracting observable information (versions, dates, URLs), followed by inferring complex semantics (attack methods, IOCs), and finally grouping packages with shared characteristics, producing structured JSON output containing the nine entity types.
(2) Entity Relationship Analysis (Fig 4b): Analyze the relationships between extracted entities and organize information centered around packages. The prompt design integrates four components: 1) task description of analyzing semantic relationships between extracted entities, 2) the intelligence content, 3) identified entities from the previous extraction step, and 4) few-shot examples as reference. The LLM analyzes the associations between entities, generating a structured JSON output.
(3) Entity Verification (Fig 4c): This stage verifies the accuracy of the extracted information through LLM-based cross-validation and self-correction. The process inputs both the original intelligence content and the structured entities from the entity relationship analysis step. The LLM performs cross-validation by comparing the extracted entities and their relationships against the original text, identifying inconsistencies or errors, and making necessary corrections to ensure output maintains accuracy and completeness.
During the prompt design phase, we performed iterative refinement based on small-scale performance observations to optimize the LLM extraction process. Based on the experimental results and analysis, we implemented two major prompt enhancements: 1) specifying a particular output structure and providing case studies to guide LLM in extracting the required information more accurately, and 2) updating the prompts with specific examples for correction. These design refinements enabled the LLM to perform the extraction task with reliable precision (as detailed in Section 4.3).
3.4. Intelligence Aggregation
Through the above methods, we can accurately extract information related to malicious packages from webpages. When multiple sources provide entity information on the same malicious package, we need to aggregate entity information from different sources to ensure comprehensive intelligence gathering. For each malicious package P, we collect a set of entity information $E=\{N,V,F,R,M,D,I,A,C,T\}$, which includes name N, version number V, discovery date F, repository URL R, attack method M, discoverer D, affected system I, attack vector A, IOC indicators C, and and collection timestamp T which represents the webpage publication time, extracted from webpage content during crawling.
Our analysis revealed that while most entity information is consistent across sources, conflicts do exist, particularly in attack methods and version numbers (5.13%), affected systems (2.45%), and repository URLs (0.41%). For these conflicting cases, we observed two patterns: (1) later webpages tend to be more accurate as they often incorporate and reference earlier findings, and (2) information confirmed by multiple sources demonstrates higher reliability.
Based on these observations, we employ two aggregation strategies: (1) direct merging for non-overlapping entity information from different sources, and (2) voting mechanism for partially overlapping information. Manual verification ensures aggregated information accuracy.
Voting Mechanism: For each package name $N$ of the malicious package information, vote on fields $V$, $R$, $M$, $D$, II, AA. After excluding $NaN$ values, count the occurrences of each value. Let $E_{i}$ be the value of entity $E$ in the ith intelligence source. Define the voting count function $count(E_{i})$ to represent the number of occurrences of value $E_{j}$. If there is a $\exists E_{i},E_{j}$ such that $count(E_{i})=count(E_{j})$, select the entry with the largest timestamp $T$: $E_{\text{final}}=\arg\max_{E_{i}}T_{i}$. For the discovery date $F$, select the earliest date: $F_{\text{final}}=\min_{i}F_{i}$. For IOCs $C$, merge all non-duplicate values: $C_{\text{final}}=\bigcup_{i}C_{i}$.
Through this voting-based aggregation mechanism, we can effectively aggregate entity information of malicious packages while ensuring the completeness of intelligence. The experimental results in  Section 4.3 validate the effectiveness of our voting-based aggregation approach, demonstrating its ability to accurately resolve conflicts and produce high-quality aggregated intelligence.
To ensure the accuracy of our final database, we implemented a two-step verification process. First, we verified package existence by validating against official npm and PyPI repository metadata APIs (npmregistry; pypistats; pypibigquery), which retain historical package information even after malicious packages are removed, ensuring extracted names correspond to actual packages rather than LLM hallucinations. Second, we confirmed maliciousness through multiple approaches: cross-referencing with established security databases (OSV, Snyk, GitHub Advisory), multi-source corroboration for packages reported by multiple independent sources, and manual verification for remaining packages, which proved efficient due to malicious packages typically being released in batches. Through this multi-layered verification approach, we established a reliable and accurate database of verified malicious packages.
4. Experiments
4.1. Research Questions
To evaluate the effectiveness of IntelliRadar, we designed our experiments around five distinct research questions:
$\bullet$ RQ1 (Effectiveness): What is the effectiveness of each key component in IntelliRadar for intelligence extraction?
$\bullet$ RQ2 (Completeness): How does IntelliRadar compare to existing databases on the completeness of collected SSC Intelligence?
$\bullet$ RQ3 (Timeliness): How does IntelliRadar compare to existing databases in terms of the timeliness of intelligence collection?
$\bullet$ RQ4 (Source Distribution): How do different intelligence sources contribute to IntelliRadar’s intelligence collection?
$\bullet$ RQ5 (Usability): How does IntelliRadar’s availability contribute to mitigating threat propagation in the downstream software supply chain ecosystem?
Our experimental setup outlines the dataset collected throughout this work. RQ1 examines the effectiveness of key components in our methodology, while RQ2 and RQ3 represent our paper’s primary contributions: the completeness and timeliness of malicious package intelligence. RQ4 investigates how different intelligence sources contribute to these core attributes. RQ5 assesses the real-world impact of our approach in protecting downstream ecosystems.
4.2. Experimental Setup
Through comprehensive data collection from 24 intelligence sources that were constructed and identified in Section 3.1, including both unstructured webpages and structured databases (OSV, Snyk, GitHub Advisory). we collected 50,586 raw webpage texts. After filtering, we retained 28,593 webpages containing security intelligence. Through LLM analysis, we found that 11,173 (39.1%) of these security-related webpages contained no malicious package names. From the remaining webpages, we extracted 35,229 potentially malicious package records. We first validated these records against official package registries, confirming that 34,982 (99.3%) packages existed in npm and PyPI registries, while 247 (0.7%) package names did not exist in these registries, indicating they were not real packages. For accuracy validation, we cross-referenced our dataset with established security databases, which verified 29,485 packages (83.7%) as malicious packages. The remaining 5,744 packages underwent our multi-source verification process, where packages reported by at least two independent sources were automatically classified as confirmed malicious (3,868 packages, representing 67.5% of the verification subset). For the remaining 1,876 single-source packages, we conducted manual verification, leveraging the observation that malicious packages are typically distributed in batches and can therefore be verified collectively when referenced in the same intelligence source. This verification identified 960 additional malicious packages, while 916 (2.6% of the original collection) were classified as benign. After this verification process, our final database contains 34,313 confirmed malicious package names (12,522 PyPI and 21,791 npm packages), achieving a precision rate of 97.4%.
4.3. RQ1: Effectiveness
Regarding the effective collection and identification of malicious package intelligence, in this section, we first investigate the contribution of each steps to effectiveness.
LLM Parameter Configuration. Throughout all experiments in this study, we maintain consistent LLM parameter settings to ensure fair comparison and reproducibility. Specifically, we configure the temperature parameter to 0 to minimize randomness in model outputs, and set top_p to 0.3 to control the diversity of generated responses. We utilize GPT-4o version gpt-4o-2024-11-20 and GPT-4o-mini version gpt-4o-mini-2024-07-18, both with knowledge cutoff of October 2023. All models are accessed through Azure-provided OpenAI API with API version 2024-02-15-preview. All models and experiments are conducted under this unified parameter configuration.
Validation of keyword text filtering. To evaluate the effectiveness of keyword text filtering, we randomly selected 250 text samples related to malicious package intelligence and 250 unrelated text samples. The experimental results shown in Table 2 indicate that the system successfully filtered 88.9% of irrelevant texts (220/250), with only 30 irrelevant samples misclassified as relevant. Meanwhile, among the 250 relevant samples, only 9 were incorrectly filtered out, resulting in a recall rate of 96.4%. To further validate our filtering effectiveness at scale, we conducted comprehensive analysis on all 50,586 original webpages using broader security-related terms. This expanded filtering identified 5,893 additional pages, from which manual examination revealed only 27 additional malicious packages (0.08% of our total collection), confirming that our original approach captured the vast majority of relevant intelligence.
Table 2. Accuracy of Keyword Filtering for Malicious Package-Related and Unrelated Texts
Metric Precision Recall F1 FN FP
Keyword Filtering 88.9% 96.4% 92.5% 9 30
Table 3. Evaluation of Different LLMs and Methods
Model Precision Recall F1 FN FP
LLaMA3.3-70B 95.63% 67.46% 79.11% 232 22
LLaMA3.1-70B 97.25% 54.56% 69.90% 324 11
Qwen2.5-72B 95.20% 75.04% 83.92% 178 27
GPT-4o-mini 91.53% 63.67% 75.10% 259 42
CTIKG (huang2024ctikg) 53.85% 12.44% 20.22% 591 72
SecBERT (host2023constructing) 0.10% 2.07% 0.20% 661 13,340
IntelliRadar few-shot 95.74% 59.89% 73.68% 286 19
IntelliRadar CoT 94.51% 91.73% 93.10% 59 38
IntelliRadar 97.91% 92.01% 94.87% 57 14
Note: Results based on the final verification stage of LLM analysis.
Table 4. Model Performance on Recent Security Reports
Model Precision Recall F1 FP FN
IntelliRadar (GPT-4o) 99.17% 88.81% 93.70% 1 15
GPT-4o-mini 87.80% 80.60% 84.05% 15 26
Qwen2.5-72B 97.96% 71.64% 82.76% 2 38
LLaMA3.1-70B 95.08% 86.57% 90.62% 6 18
LLaMA3.3-70B 92.19% 88.06% 90.08% 10 16
Validation of entity extraction. To evaluate to what extent selecting different prompting strategies, as well as different LLMs, could influence the effectiveness of IntelliRadar, we implement two sets of variants of our approach for comparison: 1) we implement our approach with different prompting strategies, IntelliRadar few-shot using only a few shots, IntelliRadar CoT implementing only CoT reasoning, to evaluate which contributes more to our complete approach (IntelliRadar) that integrates both strategies. 2) We also implement our approach with different LLMs, for instance, LLaMA3.3-70B, LLaMA3.1-70B, Qwen2.5-72B, and GPT-4o-mini, to evaluate the influence of select GPT-4o to our approach. We also compared with existing approaches: CTIKG achieved 20.22% F1-score due to lack of explicit entity definitions for malicious packages, while SecBERT performed poorly (0.20% F1) as it incorrectly identified numerous irrelevant text spans, resulting in extremely high false positives. The evaluation is performed on the 713 ground truth packages in the validation of entity extraction, treating all versions of the same package as a single entity where extraction is considered correct only when both package name and all associated versions match exactly. Our evaluation framework defines three outcomes: correctly extracted packages (TP), missed packages in webpages (FN), and incorrectly extracted package names (FP). As shown in Table 3, the experimental results demonstrate significant variations in the performance of the model in different LLMs and prompting strategies. IntelliRadar approach achieves optimal performance with a precision of 97.91%, and recall of 92.01%. The low FP count (14) indicates minimal incorrect extractions of non-existent packages, benign packages, and malformed data, while the low FN count (57) shows few malicious package entities in the text were missed during extraction, demonstrating the model’s robust ability to accurately identify malicious package information.
To address data leakage concerns, we conducted an additional evaluation on 50 security intelligence webpages published after January 1, 2025, ensuring all content postdates the training cutoff of evaluated models. As shown in Table 4, IntelliRadar achieves 93.70% F1-score with only 1 false positive, indicating strong precision on unseen data. These findings align with our evaluation results (Table 3), demonstrating that IntelliRadar consistently outperforms baseline models on both seen and unseen data. Beyond our core task of extracting package names and versions, we also evaluated other security-related entities. As shown in Table 5, our method demonstrates good performance on supplementary entities such as IOC (87.10% F1) and Discoverer (87.32% F1) extraction.
From the perspective of different prompting strategies, our results reveal that the Chain-of-Thought (CoT) approach substantially improves the recall rate from 59.89% (IntelliRadar few-shot) to 91.73% (IntelliRadar CoT). This improvement can be attributed to CoT’s systematic decomposition of the extraction task, enabling a more comprehensive identification of package entities. In particular, while IntelliRadar CoT achieves a high recall, its false positive count (38) is higher than IntelliRadar (14), demonstrating how the combination of both strategies helps maintain precision while maintaining high recall.
Table 5. Other Intelligence Extraction Performance on Security Reports
Entity Precision Recall F1
Date of Discovery 95.24% 83.33% 88.89%
Discoverer 86.11% 88.57% 87.32%
IOC 80.60% 94.74% 87.10%
Method of Attack 75.00% 78.95% 76.92%
Attack Vector 72.00% 75.00% 73.47%
Impacted Systems 64.71% 68.75% 66.67%
Repo URLs 53.85% 77.78% 63.64%
Validation of intelligence aggregation To evaluate the accuracy of our entity aggregation approach, we randomly selected 200 entities that required aggregation, which originated from 635 different intelligence source reports. Each selected entity contained conflicting information that needed to be consolidated. To prove the effectiveness of our voting mechanism, we manually label the ground truth for each of them. Our voting approach achieved an accuracy of 90.5%, successfully aggregating 181 entities correctly while incorrectly aggregating 19 entities, which confirms the majority are right in the field of malicious package intelligence, and also proves that our simple voting mechanism can fit most of conflicting cases in malicious package intelligence. This reduces the cost of manual inspection.
Response to RQ1: The text filtering effectively removes irrelevant content, with IntelliRadar outperforming all baseline models in entity extraction and relationship alignment. By combining CoT prompting with few-shot, IntelliRadar achieves an F1 score of 94.87%, surpassing other LLMs.
4.4. RQ2: Completeness
Figure 5. Comparison of Malicious Package Coverage in Different Databases
In this section, we compare the data completeness between IntelliRadar and open-source databases including Snyk, OSV and GitHub Advisory. Figure 5 provides a detailed comparative analysis result. The results show that for PyPI malicious package intelligence, only 9 entries are common across all four databases, while OSV, Snyk, and IntelliRadar share 7,007 data points. For NPM malicious package data, 11,558 entries are common across all databases, primarily because NPM malicious intelligence is reported through GitHub Advisory, and other databases collect related intelligence based on this. GitHub Advisory contains only 10 PyPI malicious packages, significantly fewer than other databases. This disparity occurs because PyPI lacks a unified reporting mechanism, and GitHub Advisory primarily focuses on formal CVE assignments rather than malicious packages. Many PyPI malicious packages are detected and removed before formal advisories are issued, creating substantial coverage gaps in centralized databases.
IntelliRadar includes data from OSV, Snyk, and GitHub Advisory to ensure comprehensive coverage, as these databases contain packages identified through internal detection tools and manual research not disclosed elsewhere. Beyond this data, our key contribution lies in extracting intelligence from other sources, identifying 17,759 NPM packages and 11,248 PyPI packages through LLM-based analysis. This includes 2,262 NPM packages and 2,566 PyPI packages not present in any existing structured databases, demonstrating our approach’s effectiveness in discovering malicious packages from dispersed sources.
The OSV database contains 17,999 NPM and 8,772 PyPI malicious packages. IntelliRadar covers all OSV database entries and identifies an additional 3,792 NPM and 3,750 PyPI malicious packages, representing 21.07% and 42.74% respectively. Analysis of these unique data sources reveals that for NPM packages, Sonatype contributes 34.83% of the intelligence, Phylum provides 19.29%, with the remaining sources including QianXin, Medium, and Socket. For PyPI packages, Medium and Sonatype contribute 24.08% and 16.26% of the intelligence respectively, with additional sources including Phylum, Checkmarx, QianXin, Reddit, and TuxCare.
Comparison with the Snyk database shows that IntelliRadar shares 13,465 NPM and 8,200 PyPI malicious packages, while identifying 8,326 additional NPM packages (38.21%) and 4,322 PyPI packages (34.52%) through analyzing open-source web pages. Of these additional NPM packages, 25.51% of the intelligence comes from GitHub and 16.90% from Sonatype, with the remainder from Phylum and Socket. For all PyPI packages collected by IntelliRadar from unstructured web content, the primary sources are Medium and Sonatype. Among the unique PyPI packages (those exclusively identified by IntelliRadar), the intelligence sources are primarily Phylum, Checkmarx, and Twitter, with these packages predominantly employing Typosquatting attacks. Notably, IntelliRadar detected 749 packages on December 4, 2023, the same day they were released. The results demonstrate that LLM-based unstructured data analysis expands the coverage of existing databases.
To quantify the real-world impact of these uniquely identified malicious packages, we analyzed their downstream adoption through download statistics. We focused this analysis on PyPI packages, as PyPI officially provides comprehensive download records through ”pypi.file_downloads” in Google BigQuery, while no comparable data exists for NPM. Our analysis revealed that the 2,566 PyPI malicious packages uniquely identified by IntelliRadar were collectively downloaded 326,740 times before being detected. Among these, 1,427 packages were downloaded more than 100 times each.
Response to RQ2: IntelliRadar is more comprehensive than well-known databases like OSV and Snyk. Additionally, it collected 4,828 (PyPI: 2,566 and NPM: 2,262) exclusive pieces of intelligence from all data sources.
4.5. RQ3: Timeliness
Figure 6. Time Intervals: IntelliRadar (Excluding OSV/Snyk) vs. OSV and Snyk
To validate how IntelliRadar outperforms existing malicious package databases in timeliness, we compared the dates when IntelliRadar and these databases recorded malicious packages. To ensure a fair comparison, we specifically evaluated malicious packages collected by IntelliRadar from sources other than OSV and Snyk databases, comparing their detection timestamps against the same packages recorded in OSV and Snyk.
Compared to the Snyk database. For 10,314 NPM malicious packages, IntelliRadar recorded the intelligence earlier than Snyk, accounting for 76.6%; among these, 6,598 packages were recorded 100.3 days earlier, with 49% of packages recorded a year in advance. For PyPI malicious packages, 5,765 packages were recorded earlier than Snyk, accounting for 70.3%, with 5,196 packages recorded more than a week earlier. Additionally, 566 malicious packages had consistent recording times, and only 189 packages (2.3%) were recorded earlier by Snyk. The earlier detection capabilities of IntelliRadar stem from our comprehensive coverage of social media and unstructured websites from security companies, enabling us to discover the latest malicious packages immediately. For the packages that Snyk detected earlier, their internal proprietary detection tools identified these specific threats before public disclosure. These delays in Snyk are attributed to its dependency on established databases, GitHub security advisories, and manual research processes (snykvulndb).
Figure 7. Distribution of Collected Malicious Intelligence Across Various Intelligence Sources
Compared to the OSV database. IntelliRadar and OSV have consistent recording times for 15,479 NPM packages, accounting for 86.0% of the total. Additionally, IntelliRadar recorded 171 malicious packages earlier than OSV, primarily collected from Sonatype and Securityaffairs (securityaffairs_blog_ref). For PyPI malicious packages, IntelliRadar recorded 4,711 earlier than OSV, with most differences within a week; 53.7% were just one day earlier, mainly from Phylum. 511 packages were recorded on the same day by both IntelliRadar and OSV. Furthermore, we extracted PyPI package release times from Google’s bigquery-public-data and NPM package release times from the NPM Registry. Analysis shows that 4,533 NPM and 549 PyPI malicious package names were identified by IntelliRadar on the day of their release, indicating our method can swiftly collect relevant intelligence once malicious packages are publicly reported. OSV’s delays are attributed to its aggregation model, which depends on multiple upstream databases (osvdata) and integrated detection tools (packageanalysis).
To demonstrate the significance of timeliness of be cognitive to malicious package intelligence, we compared the time intervals of malicious package inclusion across different databases and the number of downloads during these intervals. Figure 6 shows the distribution of time differences between IntelliRadar-OSV and IntelliRadar-Snyk for malicious packages. Notably, 4,916 malicious packages were recorded in the Snyk database on 07/03/2023, while IntelliRadar identified these package names on 24/02/2023, through the Phylum intelligence source. Google Cloud data reveals that during these 11 days, these packages were downloaded 201,556 times, averaging 41 downloads per package. Further analysis indicates an uneven distribution of downloads, with the United States accounting for 49.2% and China for 19.4%. Among these, 107 malicious packages were downloaded 100 times each, with one package named studypong downloaded 233 times. These data suggest that these packages were widely distributed and affected numerous users before being recorded by mainstream security databases. Our method enables early identification and warning before malicious packages become widespread, significantly reducing their lifecycle and minimizing their impact while protecting user safety.
Response to RQ3: 73.9% of IntelliRadar’s intelligence is recorded earlier than the OSV database, and 57.9% of PyPI intelligence is recorded earlier than the Snyk database, effectively shortening the lifecycle of malicious packages.
4.6. RQ4: Source Distribution
We conducted a quantitative analysis of intelligence sources for malicious packages in the PyPI and NPM ecosystems. As shown in the figure, the PyPI ecosystem exhibits a diversified distribution pattern, with Phylum platform (34.8%), Medium community (13.4%), and OSV database (10.04%) serving as the primary intelligence sources; in contrast, the NPM ecosystem demonstrates a highly centralized distribution, dominated by GitHub Advisory (45.85%), Sonatype (17.56%), and Phylum (8.55%). Through in-depth analysis, we found that 32.4% of malicious packages were documented by multiple data sources, while in the NPM ecosystem, 87.27% of malicious packages were recorded by only a single source, primarily attributable to NPM’s adoption of GitHub Advisory as the main channel for publishing malicious package intelligence. For packages uniquely identified by IntelliRadar, in the NPM ecosystem, these were primarily contributed by Phylum (23.61%), GitHub (21.32%), and Sonatype (21.28%); PyPI malicious packages came from Sonatype (32.50%), Medium (22.22%), Qianxin (11.1%), and Reddit, which provided 157 packages (8.77%). Regarding intelligence timeliness, our data revealed significant differences among sources. In the PyPI ecosystem, Checkmarx and Phylum led in malicious package reporting times, accounting for 44.42% and 23.25% of reports, respectively. Specifically, Checkmarx issued alerts 3.02 days earlier on average, while Phylum reported 2.77 days earlier. Similarly, in the NPM ecosystem, for packages documented by multiple sources, GitHub and Phylum demonstrated superior timeliness, accounting for 40.35% and 17.67% of early reports. GitHub Advisory identified threats 15.29 days earlier on average compared to other intelligence sources, while Phylum reported 6.62 days earlier. These time differences reflect variations in threat identification capabilities and information dissemination mechanisms among intelligence sources, providing valuable reference for building efficient security response systems.
Response to RQ4: Key intelligence sources differ between ecosystems - PyPI mainly relies on Phylum, Medium and OSV while NPM centralizes around GitHub Advisory. For timeliness, Checkmarx and Phylum lead in PyPI (reporting 3.02 and 2.77 days earlier), while GitHub dominates NPM Intelligence.
4.7. RQ5: Usability
Table 6. Comparison of IntelliRadar Downstream Mirror Retention and Inclusion Times in Other Databases
Package Name Versions PyPI Mirrors Database Downloads
Tsinghua Tencent Aliyun Douban BFSU OSV Snyk GitHub Advisory IntelliRadar (Our)
1inch 8.6 - 8.9 - - - 10/10/22 (+3) 07/10/22 110
libcontroltoolver 4.86 - - - 26/02/23 (+0) 07/03/23 (+9) 26/02/23 62
matplotlyib 1.0.0 - - 29/03/23 (+31) 26/02/23 64
pipcryptov4 1.0.0 - - - 25/06/24 (+266) 05/10/23 (+2) 03/10/23 301
libideeee 1.0.0 - - - 25/06/24 (+294) 07/09/23 (+2) 05/09/23 74
ethereum2 2.8.4, 2.8.6, 2.8.9 - - - 17/12/23 (+436) 07/10/22 331
gkjzjh146 1.3 - - - 14/05/23 (+226) 30/09/22 638
httprequesthub 2.31.0 - 2.31.4 - - - 25/06/24 (+186) 25/12/23 (+3) 22/12/23 496
logic2 0.1.4 - - - - 29/03/23 (+104) 15/12/22 520
pipsqlite3liberyV2 1.1.0 - - - - - 22/05/23 (+104) 15/05/23 145
flak7 4.5.2 - - - - 07/09/22 (+5) 02/09/22 18
simpeljson 4.5.2 - - - - 11/02/23 20/06/23 (+129) 11/02/23 67
pyward 3.0 - - - - 08/09/23 (+1) 07/09/23 559
studypong 5.66, 7.16, 8.22, 10.45 - - - - 25/02/23 (+1) 07/03/23 (+11) 24/02/23 264
reqkests 2.28.1 - - - - 11/12/22 (+2) 09/12/22 143
beautiflulsoup 1.0.0 - - - 29/03/24 (+1) 28/03/24 76
pycryptdome 4.4.2 - - - - - 25/08/22 63
1337z 4.4.7 - - - - 31/08/22 176
urllib7 1.26.12 - - - - 15/12/22 69
urllib12 1.26.12, 1.30.0 - - - - 15/12/22 296
The symbol indicates the presence of the malicious package in the mirror, while ’-’ indicates its absence. The symbol indicates missing intelligence. The notation (+n) shows the number of days the package was included later than in our IntelliRadar database. Downloads indicate the number of times the malicious package has been downloaded.
In the context of package management ecosystems, the rapid propagation of malicious packages from primary repositories to downstream mirrors presents a significant security challenge. When a malicious contributor uploads a malicious package to pypi.org, it swiftly synchronizes across various mirror sources. Due to the varied synchronization mechanisms (full or incremental) employed by different mirrors, downstream mirrors receive no formal notification when malicious packages are removed from pypi.org, potentially resulting in the prolonged persistence of these threats within mirror repositories.
Figure 8. Monthly Analysis of Supply Chain Security Intelligence Sources and Malicious Packages
To this end, we scanned the main downstream PyPI mirrors (guo2023empirical) using the intelligence database. As of 01/05/2025, we discovered 1,981 malicious packages across various PyPI mirrors, specifically 781 in the Tsinghua mirror, 548 in the Tencent mirror, 395 in the Douban mirror, 21 in the BFSU mirror, 132 in the Huawei mirror, and 104 in the Aliyun mirror. Among these malicious packages, 54.7% were uniquely identified in our database. These malicious packages had been downloaded over 86,240 times cumulatively, with the most severe case being the ethereum2 package, which had been downloaded 436 times before our identification and was identified 436 days earlier than other security databases such as Snyk. Table  6 presents the most impactful cases, highlighting the advantage of our tool in early identification.
In response to these findings, we sent batch emails to downstream mirror maintainers providing data on the detected malicious packages. We received confirmation replies from Tencent, Douban, and Tsinghua. By comparing the package status before and after our notification, we verified that the administrators had removed all the malicious packages. For real-time deployment data, we also issue monthly reports to downstream mirrors to ensure the security of the entire ecosystem.
Beyond these practical applications, we also evaluated the financial viability of IntelliRadar through a comprehensive cost analysis. We evaluated the costs of extracting malicious package information using LLMs. The initial webpage text from 24 data sources contained over 17 million tokens. After text filtering, this was reduced to 7.07 million tokens, a reduction of 58.4%. The entire process cost \$93.8, We extracted a total of 34,313 pieces of intelligence. The average cost per piece of intelligence was \$0.003. This approach proved cost-effective, and we anticipate further cost reductions with future updates from OpenAI and open-source LLMs.
As shown in Figure 8, we conducted monthly distribution analysis of intelligence sources and malicious packages. IntelliRadar filters out 130 critical intelligence pages from 2,371 original webpages monthly, identifying an average of 650 malicious packages per month from PyPI and NPM platforms. With monthly operational costs of only 7 USD, the system demonstrates exceptional cost-effectiveness. The data reveals a surge in malicious packages since 2022, indicating escalating software supply chain threats.
Response to RQ5: IntelliRadar identified and confirmed 1,981 malicious packages in downstream mirrors, with an intelligence cost of only \$0.003. Continuous monitoring for one month costs just \$7. As open-source LLMs continue to improve, these costs are expected to decrease further.
5. Discussion
$\bullet$ Intelligence Collection Beyond PyPI and NPM. In this study, we successfully collected a large amount of malicious package intelligence related to PyPI and NPM, as well as 928 other types of threat intelligence, covering component vulnerabilities, CVE vulnerabilities, and malicious package information from other package managers. It is worth noting that we discovered malicious components Prettier in the VSCode Marketplace (prettier_vscode), which uses typesquatting for attacks (prettier_vscode). We also identified the XZ Utils software supply chain attack case. Our approach can quickly adapt to the intelligence requirements of various package manager platforms and effectively expand to new security domains, providing the security community with a more timely automated intelligence discovery.
$\bullet$ Data Poisoning Attacks. To combat potential data poisoning attacks, our framework implements multiple defensive mechanisms across the intelligence processing pipeline. During source identification, we meticulously screen and validate intelligence sources to ensure strong correlation with package management ecosystems, mitigating the risk of contaminated data. The intelligence aggregation mechanism employs a voting system to cross-validate entities from multiple independent sources, minimizing the impact of poisoned data. The intelligence extraction process strengthens defense through a three-step CoT approach (entity extraction, relationship alignment, and verification) that identifies and filters inconsistent or suspicious intelligence. The effectiveness is validated by experimental results: among 34,313 pieces of intelligence collected, 29,485 (85.9%) were corroborated against authoritative security databases (OSV, Snyk, GitHub Advisory). This high correspondence rate demonstrates that our multilayered defense strategy successfully maintains the integrity and reliability of collected intelligence while effectively countering data poisoning attempts.
6. Related Works
6.1. Open Source Intelligence Analysis
Open-source intelligence (OSINT) collects information through public channels, widely used in cybersecurity. Researchers have developed automated systems to gather threat intelligence from various sources, including the internet (mulwad2011extracting), social media (sabottke2015vulnerability), developer communities (purba2023extracting), security forums, and code repositories (neil2018mining; bouwman2022helping; huang2021hackerrank). Recent work has applied LLMs for security intelligence analysis (guo2025automated), with CTIKG (huang2024ctikg) using LLMs for knowledge graph construction from cyber threat intelligence, and Høst et al. (hostconstructing) constructing knowledge graphs from vulnerability descriptions in the NVD. In software engineering, extensive research has focused on named entity recognition (NER) and relation extraction (RE) from unstructured text sources. Studies have extracted software entities from tweets (zhang2022benchmarking), Stack Overflow posts (huang2023api; huo2022arclin), and technical documents (nguyen2023software), with recent work exploring self-evolving software intelligence management (liu2025evolaris). These approaches have been applied to API mention resolution (huang2023api), crash solutions (zhang2022benchmarking), library recognition (zhang2022benchmarking), and version incompatibility detection (zhao2023knowledge). However, malicious package intelligence extraction presents unique challenges distinct from traditional software entity recognition. IntelliRadar requires: (1) distinguishing malicious package names from benign packages mentioned in the same text (where benign packages are often the legitimate targets that malicious packages attempt to mimic), (2) identifying the specific ecosystem (npm vs PyPI) for each package, and (3) performing semantic-level understanding to extract complex attack vectors and indicators of compromise.
6.2. Software Supply Chain Security
Software supply chain security faces severe challenges, with package management systems becoming hotspots for malicious activities. Attackers inject malicious code through methods like code tampering, dependency confusion, and typesquatting, leading to data leakage and system damage. DL Vu et al. reveal various attack methods and anti-detection techniques (vu2023bad; duan2020towards; gu2023investigating). To address these threats, researchers have employed strategies ranging from static and dynamic analysis to machine learning. The MalOSS framework proposed by Duan et al. (duan2020towards) demonstrates similarities in malicious behaviors across different languages through metadata and API call sequence analysis. Huang et al. utilize graph-based behavior modeling to detect malicious packages (huang2024spiderscan). Zhang et al.(zhang2023malicious) utilize deep learning to understand malicious software’s behavioral semantics, while Liang et al.(liang2023needle) demonstrate the effectiveness of anomaly analysis algorithms. Additionally, GitHub, as the largest source code repository, may serve as a channel for distributing malicious code, with inconsistencies between distributed packages and source code potentially indicating malicious injection (vu2021lastpymile). However, existing detection methods face a critical limitation: detection results are scattered across security blogs and unstructured webpages, failing to reach downstream users effectively. Even with numerous detection tools, malicious packages persist because intelligence remains fragmented. IntelliRadar addresses this through a fundamentally different approach—systematically analyzing malicious package intelligence from dispersed sources, covering both historical and newly reported packages.
7. Conclusion
In the open-source ecosystem, package managers like PyPI and NPM lack up-to-date intelligence databases, allowing malicious packages to persist. To address this, we developed IntelliRadar, a system leveraging multi-source data collection and LLMs to gather 34,313 pieces of intelligence from 24 sources. The system enables earlier detection of malicious packages, reducing their impact, at \$0.003 per intelligence item. This work presents an efficient solution to enhance open-source software supply chain security.
Acknowledgment
This work was supported by the National Key Research and Development Program of China (Grant No. 2024YFF0908000). This research is supported by the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG4-GC-2023-008-1B); by the National Research Foundation Singapore and the Cyber Security Agency under the National Cybersecurity R&D Programme (NCRP25-P04-TAICeN); and by the Prime Minister’s Office, Singapore under the Campus for Research Excellence and Technological Enterprise (CREATE) Programme. Any opinions, findings and conclusions, or recommendations expressed in these materials are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore, Cyber Security Agency of Singapore, Singapore.